OpenStudy (butterflyhope):

Simplify: 3[x + 2(x - 3)]. Show work. Please help, thanks!

OpenStudy (anonymous):

ok im going to help you work it out ok?

OpenStudy (butterflyhope):

Ok ^-^

OpenStudy (anonymous):

gimme a sec to work it out myself

OpenStudy (butterflyhope):

Kk

OpenStudy (anonymous):

ok lets work inside the brackets first. Do you know what you have to do first?

OpenStudy (anonymous):

you're going to take the number in front of the parenthesis and do the foil method. so 2*x and 2*-3

OpenStudy (anonymous):

so then you get 3[x+2x-6)]

OpenStudy (butterflyhope):

Ok I get it so far :p

OpenStudy (anonymous):

next combine like terms 3[3x-6]

OpenStudy (anonymous):

add x=2x and you get 3x

OpenStudy (anonymous):

imean x+2x=3x

OpenStudy (butterflyhope):

Yea :)

OpenStudy (anonymous):

now finally take the number OUTSIDE the brackets and distribute. Now what do you get?

OpenStudy (butterflyhope):

3 x 3x = 9x. 3 x -6 = -18. 9x times -18 = -162x? I'm tired and delusional, so please bear with me if I get it wrong 99 times. xD

OpenStudy (anonymous):

lol its ok i understand, ive been at school ALL day lol, but you wikk get 9x-18 3*3x=9x 3*-6=-18

OpenStudy (butterflyhope):

Thanks :)

OpenStudy (anonymous):

NP :)